The PMC was unlocked when I came in ~10mins ago. The wall StripTool traces suggest it has been this way for > 8hours. I was unable to get the PMC to re-lock by using the PMC MEDM screen. The c1psl slow machine responded to ping, and I could also telnet into it. But despite burt-restoring c1psl, I could not get the PMC to lock. So I re-started c1psl by keying the crate, and then burt-restored the EPICS values again. This seems to have done the trick. Both the PMC and IMC are now locked.
Unrelated to this work: It looks like some/all of the FE models were re-started. The x3 gain on the coil outputs of the 2 ITMs and BS, which I had manually engaged when I re-aligned the IFO on Monday, were off, and in general, the IMC and IFO alignment seem much worse now than it was yesterday. I will do the re-alignment later as I'm not planning to use the IFO today.